光情報システム工学論 2 1 単位 (選択)

Lecture in Optical Information Systems, Part 2  仁木登 · 教授/ システム創生工学専攻光システム工学コース光情報システム工学講座

【授業目的】イメージング技術，知的画像処理に関する最近の研究動向，今後の

展望について理解することを目的とする．【授業概要】イメージング技術，知的画像処理に関する最近の研究動向，今後の

展望について講義する．【授業形式】 [形態] 【キーワード】イメージング技術，知的画像処理【先行科目】 [先行科目] 【関連科目】『3 次元画像処理』 (0.5)，『バーチャルリアリティ技術』 (0.5) 【履修要件】 [要件] 【履修上の注意】 [注意] 【到達目標】イメージング技術，知的画像処理に関する最近の研究動向，今後の

展望について理解する．【授業計画】

1. 3 次元 X 線 CT 技術の研究動向と今後の展望 (1)

2. 3 次元 X 線 CT 技術の研究動向と今後の展望 (2)

3. 3 次元 X 線 CT 技術の研究動向と今後の展望 (3)

4. 3 次元 X 線 CT 技術の研究動向と今後の展望 (4)

5. 3 次元 X 線 CT 技術の研究動向と今後の展望 (5)

6. MR イメージング技術の研究動向と今後の展望 (1)

7. MR イメージング技術の研究動向と今後の展望 (2)

8. MR イメージング技術の研究動向と今後の展望 (3)

9. MR イメージング技術の研究動向と今後の展望 (4)

10. MR イメージング技術の研究動向と今後の展望 (5)

11. 知的画像処理の研究動向と今後の展望 (1)

12. 知的画像処理の研究動向と今後の展望 (2)

13. 知的画像処理の研究動向と今後の展望 (3)

14. 知的画像処理の研究動向と今後の展望 (4)

15. 知的画像処理の研究動向と今後の展望 (5) 【成績評価基準】レポートにより評価する．【教科書】授業中に紹介する．【参考書】 [参考資料] 【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216835 【対象学生】 [対象学生] 【連絡先】 [連絡先] 【備考】授業を受ける際には， 2 時間の授業時間毎に 2 時間の予習と 2 時間の復 習をしたうえで授業を受けることが，授業の理解と単位取得のために必要で ある．

Lecture in Optical Information Systems, Part 2 1 unit (selection)

Noboru Niki · Professor / Optical Materials and Devices, Optical Systems Engineering, Systems Innovation Engineering

Target⟩ Introduction to recent topics on imaging technologies and intelligent image

processing. Outline⟩ This course introduces recent technical information on imaging technolo-

gies and intelligent image processing and describes their future directions. Style⟩ [形態] �������⟩ imaging technologies, intelligent image processing Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ “Three-dimensional Image Processing”(0.5), “Virtual Reality”

(0.5)

Requirement⟩ [要件] Notice⟩ [注意] Goal⟩ To understand the recent topics on imaging technologies and intelligent

image processing.

Schedule⟩

1. Recent topics of three-dimensional CT technologies (1)

2. Recent topics of three-dimensional CT technologies (2)

3. Recent topics of three-dimensional CT technologies (3)

4. Recent topics of three-dimensional CT technologies (4)

5. Recent topics of three-dimensional CT technologies (5)

6. Recent topics of MR imaging technologies (1)

7. Recent topics of MR imaging technologies (2)

8. Recent topics of MR imaging technologies (3)

9. Recent topics of MR imaging technologies (4)

10. Recent topics of MR imaging technologies (5)

11. Recent topics of intelligent image processing (1)

12. Recent topics of intelligent image processing (2)

13. Recent topics of intelligent image processing (3)

14. Recent topics of intelligent image processing (4)

15. Recent topics of intelligent image processing (5) Evaluation Criteria⟩ Report 100% Textbook⟩ Reference books are introduced to each topic. Reference⟩ [参考資料] Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216835 Student⟩ [対象学生] Contact⟩ [連絡先] Note⟩ Preparation (2hrs) and Review (2hrs) are required to take this lecture (2hrs).